# MPCM-Net: Multi-scale network integrates partial attention convolution with Mamba for ground-based cloud image segmentation

### v1

@@ II-C Ground-based Cloud Image Dataset
With the continuous advancement of DL techniques in ground-based cloud image segmentation, all-sky imagers have emerged as the instrument of choice for many research institutions and centers to capture high-resolution sky imagery. To this end, these organizations have begun to release high-quality, publicly available ground-based cloud image datasets with the aim of further improving the accuracy of cloud segmentation.
@@
Li et al. introduced the UTILITY dataset for analyzing statistical characteristics of clouds and sky elements [dataset1]. They manually converted 32 of these images into binary mask images with a resolution of 682 $\times$ 512 using a Voronoi polygon-based region generator. Dev et al. made publicly available a large-scale annotated database of sky/cloud images to the research community, named the Singapore whole sky imaging segmentation database (SWIMSEG) [dataset8]. The SWIMSEG dataset comprises 1013 images captured by the wide angle high-resolution sky imaging system (WAHRSIS), developed and deployed at Nanyang Technological University, Singapore, with each image having a resolution of 600 $\times$ 600 pixels [dataset3]. Additionally, Dev et al. released the Singapore whole sky nighttime imaging segmentation database (SWINSEG) [dataset2], a dedicated dataset for nighttime cloud segmentation, consisting of 115 images at a resolution of 500 $\times$ 500 pixels. Fabel et al. collected data using an all-sky imager based on a commercial surveillance camera (Mobotix model Q25), assembling a total of 770 all-sky images (ASI) with a resolution of 512 $\times$ 512 pixels [dataset4]. Building upon the SWIMSEG, Park et al. incorporated additional hazy and overcast images captured by a Waggle sensor node equipped with a camera, all of which have a resolution of 300 $\times$ 300 pixels [dataset5]. Liu et al. published the Tianjin Normal University large-scale cloud detection database (TLCDD), which contains 5000 cloud images. All images are stored in PNG format with a pixel resolution of 512 $\times$ 512 [dataset6]. Kalisch et al. employed an all-sky imager developed by the Leibniz Institute of Marine Sciences (IFM-GEOMAR) at the University of Kiel. The system is equipped with a digital CCD camera featuring a fisheye lens, providing a 183° field of view. The sky images are stored in 30-bit color JPEG format with a maximum resolution of 3648 $\times$ 2736 pixels [dataset7]. Shi et al. selected 300 images from the dataset released by [dataset7], cropped them to a resolution of 1704 $\times$ 1704, and publicly released the ground-based cloud image fine-grained segmentation dataset. By incorporating radiation characteristics, this dataset supports fine-grained segmentation of ground-based cloud images into five categories: cumulonimbus and laminatus, cumulus, stratus, cirrus, and sky[CloudFU-Net].
@@
Despite the contributions these datasets have made to the development of the community, they overlook the issue of blurred cloud pixel boundaries around radiative sources, which arises due to solar-induced color distortion effects on cloud formations.
@@
Output : Segmented image $Y_{i}\in\{0,1\}^{H\times W}$
@@ 19 $Y\leftarrow\text{Conv}_{1\times 1}(\text{Concat}(f_{s},u_{1}))$
20 $\mathcal{L}\leftarrow\alpha L_{focal}+\beta L_{dice}$
@@ III Proposed Method
In this section, we first formulate the problem of ground-based cloud image segmentation and then introduce the overall design and core idea of the proposed MPCM-Net, followed by a detailed description of its key components.
@@ III-A Problem Description
The task of ground-based cloud image segmentation can be formally defined as follows. Given an input ground-based cloud image $X_{i}\in\mathbb{R}^{3\times H\times W}$, the goal is to generate a corresponding segmentation mask $Y_{i}\in\{0,1\}^{H\times W}$, where each pixel is classified as either cloud or non-cloud. Such imagery captures high-resolution cloud structures and dynamic evolutionary patterns, which inherently exhibit multi-scale variations in both texture and shape, as illustrated in Fig. 1 (a). Moreover, as shown in Fig. 1 (b), cloud boundaries, especially at different scales and near radiation sources, often appear translucent and are visually distinct from the main cloud bodies. These characteristics cause conventional convolution-based models to lose fine local features during encoding, leading to inaccurate boundary segmentation and further causing “ghosting effects” in subsequent cloud motion extrapolation tasks. Therefore, the core challenge lies in designing a segmentation model that can effectively extract and integrate multi-scale contextual features while preserving fine boundary details. Motivated by this, we propose a novel network architecture that incorporates partial convolution and a Mamba-based decoder to enhance both segmentation accuracy and inference efficiency.
@@ III-B Overview of Structure
The core idea of our method is to explicitly model multi-scale cloud features and translucent boundaries through a dedicated encoder-decoder architecture. The encoder leverages partial convolutions to enhance multi-scale feature interactions, thereby effectively capturing characteristics of cloud structures across varying scales. Simultaneously, the decoder employs a Mamba-based linear-complexity structure to restore spatial and semantic details, enabling collaborative boundary features learning from both spatial and semantic domains.
The proposed MPCM-Net, as illustrated in Fig. 2, adopts a classic encoder-decoder architecture. The procedure of our MPCM-Net is described in Algorithm 1. Within the encoder stage, we introduce the Multi-scale Partial Attention Convolution Encoding (MPAC) to enhance feature interaction through partial convolutions while adaptively extracting multi-scale contextual structural information from ground-based cloud imagery. The MPAC comprises two key components: the Multi-scale Partial Convolution (MPC) block, which includes Partial Channel Module (ParCM) and Partial Spatial Module (ParSM), as shown in Fig. 3 (a), as well as the Multi-scale Partial Attention (MPA) block, which includes Partial Attention Module (ParAM) and ParSM, as shown in Fig. 3 (e). Within this framework, the ParCM leverages partial channel information and channel attention to enable global spatial interaction across multi-scale cloud formations. The ParSM efficiently blends channel information using 1 $\times$ 1 convolutions and spatial attention mechanisms. The ParAM employs a novel attention mechanism to expand the receptive field of the model, strengthening multi-scale feature perception capabilities for cloud structures. In the decoder stage, to mitigate boundary information loss during cloud segmentation, we design a Multi-scale Mamba Decoding (MMD) incorporating Multi-scale Mamba Block (M2B). Crucially, the M2B integrates a Spatial-Semantic Hybrid Domain (SSHD) mechanism, enabling comprehensive feature learning across global and local scales while maintaining linear computational complexity. Collectively, the MPCM-Net encoder extracts robust multi-scale contextual features through synergistic partial channel and spatial information processing. Subsequent feature enhancement by the MMD across spatial and cross-scale dimensions ultimately yields precise segmentation results suitable for multi-scale representation of ground-based cloud imagery. The following subsections elaborate on the details of each module.
@@ III-C Multi-scale Partial Attention Convolution Encoding
Effectively learning multi-scale contextual information from ground-based cloud imagery under complex weather conditions is crucial for accurate cloud segmentation. This capability enables the extraction of multi-scale features representing the same cloud cluster across consecutive frames. Dilated convolution methods have proven effective for multi-scale feature extraction by applying kernels of varying sizes to the input.
However, most existing multi-scale dilated pyramid approaches simply fuse features across scales through channel concatenation, neglecting the partial feature validity and interoperability in both spatial and channel dimensions, thereby impairing global information interaction. Moreover, when employing attention mechanisms, prevalent encoders often disregard the computational overhead and latency impacts on model efficiency. This oversight prevents most ground-based cloud segmentation methods from achieving an optimal balance between segmentation accuracy and inference speed.
To this end, we introduce partial convolution mechanisms to enhance the extraction of potential feature representations within distinct local regions for multi-scale cloud segmentation. Concurrently, inspired by MALA [MALA], a novel approach that synergistically combines the precision of softmax attention with the low complexity of linear attention. This strategy enables our model to achieve an optimal trade-off between acquiring local feature information with global interaction capabilities and maintaining computationally efficient floating-point operations (FLOPs).
As depicted in the left side of Fig. 2, our encoder architecture comprises four consecutive convolutional layers. The first three layers consist of MPC blocks, while the last layer utilizes the MPA block. Each of the layers precedes a Multi-scale Extraction Layer (MEL), as shown in Fig. 3 (a). The MEL serves for multi-scale feature extraction. Specifically, the input features $F\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ are channel, height, and width dimensions, are going through three strip convolutions of different scales after the $1\times 1$ convolution (the kernel size and stride are all 1), batch normalization (BN), and rectified linear unit (ReLU). Here, the scales of the strip convolutions are 7, 11, and 21, respectively. The reason why we chose the strip convolution is that the cloud clusters are scale varieties, while also reducing the computational complexity of the model. Notice that we employ the same strategy to set the size of convolution kernels as same as the literature [SegNeXt, LG-Umer].
In contrast to the conventional ground-based cloud image segmentation encoders, our MPAC addresses the challenge of scale variations in the same cloud cluster across different frames. This is achieved by adaptively extracting multi-scale contextual information through a combination of multi-scale strip dilated convolutions and partial convolution mechanisms. Notice that we employ the same strategy to set the layers as the literature [PartialNet] to ensure network inference capability. The encoder takes the input image to obtain feature maps of different scales through four successive stages. Local features and semantic information can be obtained with each layer from the encoder. We denote the feature map of each layer as $f_{i}(i\in[1,2,3,4])$. The size of $f_{i}$ is $2^{i-1}\cdot 64\times\left(H/2^{i}\right)\times\left(W/2^{i}\right)(i\in[1,2,3,4])$.
@@ III-C1 Partial Channel Module
To mitigate efficiency degradation caused by redundancy across multi-scale feature channels in ground-based cloud segmentation, we propose a partial channel mechanism integrated with attention operations for global information interaction. The implementation of this approach effectively balances computational efficiency and segmentation accuracy by leveraging the representational efficacy and attention of partial feature channels. Specifically, as shown in Fig. 3 (b), the input feature channels are partitioned using a channel-splitting ratio $R$, set to 1/4 as a hyperparameter, consistent with FasterNet [FasterNet]. The subset of $X_{C\times R}$ undergoes a 3 $\times$ 3 convolution to aggregate global contextual information. Concurrently, the complementary $X_{C\times(1-R)}$ leverages channel attention mechanisms to capture spatial interactions. Unlike PartialNet [PartialNet], which targets general segmentation tasks, our design addresses the unique challenges of scale variations and stochastic motion in ground-based cloud imagery. We therefore introduce coordinate attention (CA) instead of SE-Net to enhance channel-wise feature representation of cloud structures, as illustrated in Fig. 3 (d).
First, we apply global average pooling to input feature $X_{2}\in\mathbb{R}^{C\times\left(1-R\right)\times H\times W}$ along the horizontal and vertical directions as $\tilde{X}_{w}\in\mathbb{R}^{C\times\left(1-R\right)\times H\times 1}$ and $\tilde{X}_{h}\in\mathbb{R}^{C\times\times 1\times\left(1-R\right)}$, respectively. The resulting feature maps are concatenated to form descriptor $\tilde{X}\in\mathbb{R}^{C\times\left(1-R\right)\times 1\times\left(H+W\right)}$. Next, a convolutional operation (comprising 1×1 convolution, BN, and ReLU) processes $\tilde{X}$ to yield descriptor $\bar{X}$, which simultaneously encodes horizontal and vertical coordinate information. This design facilitates the extraction of features from cloud masses exhibiting stochastic motion. Subsequently, $\bar{X}$ is split along its width and height dimensions to isolate direction-specific features as $\bar{X}_{w}$ and $\bar{X}_{h}$. These features undergo separate 1 $\times$ 1 convolution followed by sigmoid activation, generating attention weights oriented along their respective axes. By incorporating directional priors, these weights are guided to capture fine-grained motion characteristics of cloud formations, while adaptive channel importance weighting is simultaneously achieved. The final output $\hat{X}$ integrates directional awareness through this process, formally described by:
$\begin{split}\tilde{X}&=\mathrm{Concat}\left(\tilde{X_{w}},\ \tilde{X_{h}}\right)\\ &=\mathrm{Concat}\left(G_{w}\left(X_{2}\right),\ G_{h}\left(X_{2}\right)\right)\end{split}$ (1)
@@
$\bar{W}_{w}$ $=\sigma\left(\mathrm{conv}\left(\bar{X}_{w}\right)\right)$ (3)
$\bar{W}_{h}$ $=\sigma\left(\mathrm{conv}\left(\bar{X}_{h}\right)\right)$ (4)
$\hat{X}$ $=X_{2}\ast\bar{W}_{w}\ast\bar{W}_{h}$ (5)
where $\mathrm{Concat}\left(\cdot\right)$ denotes the concatenation of the feature maps, $G_{w}\left(\cdot\right)$ and $G_{h}\left(\cdot\right)$ denote the global average pooling along the width dimension and height dimension, respectively. $\mathrm{Conv}\left(\cdot\right)$ denotes the convolutional operation, $\mathrm{conv}\left(\cdot\right)$ denotes the 1 $\times$ 1 convolution. $\sigma\left(\cdot\right)$ denotes the sigmoid function and $\ast$ denotes the element multiplication. The output of the ParCM is $C_{x}$:
[…]

### v2

@@ II-C Ground-based Cloud Image Dataset
The proliferation of DL techniques is predicated on the availability of high-quality datasets captured by all-sky imagers. Several public benchmarks have been released to facilitate research.
@@
Li et al. introduced the UTILITY dataset for analyzing statistical characteristics of clouds and sky elements [26]. Thirty-two images were manually converted into binary mask images with a resolution of 682 $\times$ 512 using a Voronoi polygon-based region generator. Dev et al. made publicly available a large-scale annotated database of sky/cloud images to the research community, named the Singapore whole sky imaging segmentation database (SWIMSEG) [6]. The SWIMSEG dataset comprises 1,013 images captured by the wide angle high-resolution sky imaging system (WAHRSIS), developed and deployed at Nanyang Technological University, Singapore, with each image having a resolution of 600 $\times$ 600 pixels [8]. Additionally, Dev et al. released the Singapore whole sky nighttime imaging segmentation database (SWINSEG) [42], a dedicated dataset for nighttime cloud segmentation, consisting of 115 images at a resolution of 500 $\times$ 500 pixels. Fabel et al. collected data using an all-sky imager based on a commercial surveillance camera (Mobotix model Q25), assembling a total of 770 all-sky images (ASI) with a resolution of 512 $\times$ 512 pixels [11]. Building upon the SWIMSEG, Park et al. incorporated additional hazy and overcast images captured by a Waggle sensor node equipped with a camera, all of which have a resolution of 300 $\times$ 300 pixels [35]. Liu et al. published the Tianjin Normal University large-scale cloud detection database (TLCDD), which contains 5,000 cloud images. All images are stored in PNG format with a pixel resolution of 512 $\times$ 512 [49]. Kalisch et al. employed an all-sky imager developed by the Leibniz Institute of Marine Sciences (IFM-GEOMAR) at the University of Kiel. The system is equipped with a digital CCD camera featuring a fisheye lens, providing a 183° field of view. The sky images are stored in 30-bit color JPEG format with a maximum resolution of 3648 $\times$ 2736 pixels [24]. Shi et al. selected 300 images from the dataset released by [24], cropped them to a resolution of 1704 $\times$ 1704, and publicly released the fine-grained segmentation dataset. By incorporating radiation characteristics, this dataset supports fine-grained segmentation into five categories: cumulonimbus and laminatus, cumulus, stratus, cirrus, and sky[39].
@@
While these datasets have significantly advanced the field to the development of the community, they frequently neglect the challenge of blurred cloud pixel boundaries around radiative sources, which arises due to solar-induced color distortion effects on cloud formations.
Figure 1: (a) illustrates multi-scale clouds. The red, yellow, and green solid line rectangles represent large, medium, and small-scale cloud clusters, respectively. Across sequential frames, identical cloud formations exhibit varying scale variations. (b) illustrates the examples of elongated strip-like cloud bands near the horizon in CSRC. The yellow and red dashed rectangles indicate the zenith region (image center) and the horizon region (image edge), respectively. (c) illustrates the incomplete extraction of the boundary of clouds. Note that the results are from U-Net.
@@
Output : Segmented image $Y_{i}\in\{0,1,2,3\}^{H\times W}$
@@ 19 $Y\leftarrow\text{Conv}_{1\times 1}(\text{Concat}(f_{s},u_{1}))$
20 $\mathcal{L}\leftarrow\lambda_{1}L_{focal}+\lambda_{2}L_{dice}$
@@ III Proposed Method
This section mathematically formalizes the ground-based cloud image segmentation task and subsequently delineates the architectural principles of the proposed MPCM-Net.
@@ III-A Problem Description
Formally, given an input image $X_{i}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ are channel, height, and width dimensions, respectively, the goal is to generate a corresponding segmentation mask $Y_{i}\in\{0,1,2,3\}^{H\times W}$, where each pixel is classified as either cloud or non-cloud. Ground-based imagery inherently captures high-resolution cloud structures characterized by dynamic evolutionary patterns and pronounced multi-scale variations in both texture and morphology, as illustrated in Fig. 1 (a). Furthermore, as depicted in Fig. 1 (b), cloud boundaries, especially at different scales and near radiation sources, often manifest as translucent regions. These regions are spectrally distinct from the opaque cloud core yet lack sharp gradients against the sky background. These characteristics cause conventional convolution-based models to lose fine local features during encoding, leading to inaccurate boundary segmentation and further causing “ghosting effects” during cloud motion extrapolation tasks. Consequently, the primary challenge lies in designing a segmentation model that can effectively extract and integrate multi-scale contextual features while preserving fine boundary details. Motivated by this, we propose a novel network architecture that incorporates partial convolution and a Mamba-based decoder to enhance both segmentation accuracy and inference efficiency.
@@ III-B Overview of Structure
The MPCM-Net adopts a classic encoder-decoder paradigm (Fig. 2) designed to explicitly model multi-scale cloud features and resolve boundary ambiguity through a hierarchical feature interaction mechanism. The architecture comprises two primary subsystems: the MPAC encoder and the Multi-scale MMD module.
Algorithm 1 details the procedural execution of the network. In the encoding phase, the MPAC extracts robust contextual features via a synergistic integration of the MPC block and the MPA block. The MPC further decomposes into the ParCM and the ParSM (Fig. 3), ensuring efficient feature extraction across dimensions.
Subsequently, the decoder employs the MMD, which incorporates a M2B to reconstruct spatial details. A critical innovation within the M2B is the SSHD mechanism, which facilitates comprehensive feature learning across both global and local receptive fields by leveraging the linear complexity of State Space Models. The following subsections provide a granular explication of these components.
@@ III-C Multi-scale Partial Attention Convolution Encoding
Existing methodologies often extract multi-scale information from cloud clusters across consecutive frames, a process that frequently relies on standard dilated convolutions. This reliance typically neglects inter-channel feature interoperability and incurs significant latency penalties. To mitigate these limitations, the MPAC module introduces a partial convolution mechanism [12] designed to enhance feature representation within distinct local regions.
As illustrated in the left panel of Fig. 2, the encoder comprises four consecutive convolutional stages. The initial three stages utilize Multi-scale Partial Convolution (MPC) blocks, while the final stage employs the Multi-scale Partial Attention (MPA) block. Each stage is preceded by a Multi-scale Extraction Layer (MEL), as shown in Fig.3 (a). The MEL transforms input features $F\in\mathbb{R}^{C\times H\times W}$ through a CBR block comprising a $1\times 1$ convolution, Batch Normalization (BN), and a Rectified Linear Unit (ReLU) activation. Subsequently, the features traverse three parallel strip convolutions with kernel scales of 7, 11, and 21, respectively. This kernel sizing strategy aligns with established protocols in segmentation literature [16, 34]. The encoder generates a hierarchical feature pyramid $f_{i}$ for $i\in\{1,2,3,4\}$, where the resolution of $f_{i}$ is $H/2^{i}\times W/2^{i}$ with channel dimensions scaling as $64\cdot 2^{i-1}$.
Figure 3: The structure of (a) the MPC block and (e) the MPA block; (b), (c), and (g) are the structures of the ParCM, the ParSM, and the ParAM, respectively; (d) is the structure of the CA.
We incorporate the strip convolution mechanism to effectively capture the anisotropic characteristics of cloud formations[20, 9]. Ground-based imagers project the hemispherical sky onto a 2D plane, causing cloud bands to appear as elongated “strips”, particularly near the horizon[13]. To provide an intuitive justification for the necessity of strip convolution, Fig. 1 (b) illustrates the pronounced differences in cloud morphology between the zenith and horizon regions in CSRC. Near the image center, cloud formations preserve their complete geometric characteristics. Toward the image periphery, however, the mirroring-induced compression inherent to spherical projection deforms the clouds into banded, strip-like structures. Conventional square kernels ($k\times k$) introduce redundancy and background noise when processing such geometries. By decomposing a $k\times k$ convolution into a horizontal ($1\times k$) and a vertical ($k\times 1$) strip convolution, the operation can be formulated as: $Y=f(X\cdot W_{1\times k}+X\cdot W_{k\times 1})$, where $\cdot$ denotes the convolution operation. This decomposition reduces computational complexity from $O(k^{2})$ to $O(2k)$, a critical optimization for enabling real-time inference in onsite PV power forecasting systems.
@@ III-C1 Partial Channel Module
To alleviate the computational redundancy inherent in multi-scale feature channels, we propose a partial channel mechanism integrated with coordinate attention. As depicted in Fig. 3 (b), given a feature map $X\in\mathbb{R}^{C\times H\times W}$, input feature channels are partitioned based on a splitting ratio $r$ (empirically set to 0.25 based on the trade-off between FLOPs reduction and feature redundancy, consistent with FasterNet [3]). The channel dimension is split into a primary part $C_{prim}$ and a secondary part $C_{sec}$ as follows: $C_{prim}=\lfloor C\cdot r\rfloor$ and $C_{sec}=C-C_{prim}$. Accordingly, the input has been split into two parts: $X_{prim}\in\mathbb{R}^{C_{prim}\times H\times W}$ and $X_{sec}\in\mathbb{R}^{C_{sec}\times H\times W}$. The subset $X_{prim}$ is processed via a $3\times 3$ convolution to aggregate global context, while the complementary sub-feature $X_{sec}$ leverages a channel attention mechanism to model spatial interactions.
Unlike PartialNet [23], which targets general segmentation tasks, our design addresses the unique challenges of scale variations and stochastic motion. We therefore introduce coordinate attention (CA) instead of SE-Net[21] to enhance channel-wise feature representation of cloud structures, as illustrated in Fig. 3 (d).
Specifically, global average pooling is applied to the input $X_{sec}$ along the horizontal and vertical axes, yielding coordinate vectors $\tilde{X}_{w}\in\mathbb{R}^{C_{sec}\times H\times 1}$ and $\tilde{X}_{h}\in\mathbb{R}^{C_{sec}\times 1\times W}$. These vectors are concatenated to form a spatial descriptor $\tilde{X}\in\mathbb{R}^{C_{sec}\times 1\times(H+W)}$, which encodes directional information essential for tracking cloud motion. This descriptor undergoes a convolutional transformation to produce $\bar{X}\in\mathbb{R}^{C_{sec}\times 1\times(H+W)}$, which is subsequently split into direction-specific tensors $\bar{X}_{w}\in\mathbb{R}^{C_{sec}\times H\times 1}$ and $\bar{X}_{h}\in\mathbb{R}^{C_{sec}\times 1\times W}$. These tensors are processed via sigmoid activation to generate attention weights $\bar{W}_{w}\in\mathbb{R}^{C_{sec}\times H\times 1}$ and $\bar{W}_{h}\in\mathbb{R}^{C_{sec}\times 1\times W}$. This process injects directional priors into the feature map, allowing the model to focus on the fine-grained motion characteristics of cloud boundaries. The final output $\hat{X}$ is obtained via:
$\begin{split}\tilde{X}&=\left[\tilde{X_{w}},\ \tilde{X_{h}}\right]\\ &=\left[G_{w}\left(X_{sec}\right),\ G_{h}\left(X_{sec}\right)\right]\end{split}$ (1)
@@
$\bar{W}_{w}=\sigma\left(\mathrm{conv}\left(\bar{X}_{w}\right)\right),\bar{W}_{h}=\sigma\left(\mathrm{conv}\left(\bar{X}_{h}\right)\right)$ (3)
$\hat{X}=X_{2}\odot\bar{W}_{w}\odot\bar{W}_{h}$ (4)
where $\left[\cdot,\cdot\right]$ denotes the concatenation, $G_{w}\left(\cdot\right)$ and $G_{h}\left(\cdot\right)$ denote the global average pooling along the width dimension and height dimension, respectively. $\mathrm{Conv}\left(\cdot\right)$ denotes the convolutional operation, $\mathrm{conv}\left(\cdot\right)$ denotes the 1 $\times$ 1 convolution. $\sigma\left(\cdot\right)$ denotes the sigmoid function and $\odot$ denotes the element-wise multiplication. The output of the ParCM is $I_{x}$:
$I_{x}=\left[\mathrm{conv_{3\times 3}}\left(X_{prim}\right),\hat{X}\right$] (5)